Calculate (in MeV) the total binding energy and the binding energy per nucleon (a) for 3H and (b) for 3He.

Everything I've tried is wrong...
Answer #1

a)

For Tritum (3H)

No of protons = 1

No of electrons = 1

No of neutrons = 2

So, total mass of 1 proton + 1 electron + 2 neutron =

1*(1.67493*10^-27) + 1*(9.10938*10^-31) + 2*(1.672622*10^-27) = 5.021065*10^-27 kg

But mass of 3H = 3.016049 amu =  3.016049*1.6605389*10^-27 = 5.008267*10^-27 kg

So, difference in mass = 5.021065*10^-27 - 5.008267*10^-27 = 1.2798*10^-29 kg

So, Total Binding energy = 1.2798*10^-29*(3*10^8)^2

= 1.15*10^-12 J

= 1.15*10^-12/(1.6*10^-19) = 7.19*10^6 eV = 7.19 MeV <------answer

b)

For 3He,

No of protons = 2

No of electrons = 2

No of neutrons = 1

So, total mass of 1 proton + 1 electron + 2 neutron =

2*(1.67493*10^-27) + 2*(9.10938*10^-31) + 1*(1.672622*10^-27) = 5.024304*10^-27 kg

But mass of 3He = 3.0160293 amu =  3.0160293*1.6605389*10^-27 = 5.008234*10^-27 kg

So, difference in mass = 5.024304*10^-27 - 5.008234*10^-27 = 1.607*10^-29 kg

So, Total Binding energy = 1.607*10^-29*(3*10^8)^2

= 1.4463*10^-12 J

= 1.4463*10^-12/(1.6*10^-19) = 9.04*10^6 eV = 9.04 MeV <------answer